OpenStudy (anonymous):

The number of atoms of platinum (Pt) in the universe is estimated to be 1.4 times the number of atoms of silicon (Si) in the universe. The number of atoms of silver (Ag) in the universe is estimated to be 0.33 times the number of atoms of silicon in the universe. Based on these estimates, the number of atoms of platinum in the universe is approximately how many times the number of atoms of silver in the universe? Can anyone show me how to solve this, please?

OpenStudy (inkyvoyd):

*you'll want to use three variables in this case. X for Pt, y for Si, and z for Ag

OpenStudy (anonymous):

Okay

OpenStudy (inkyvoyd):

now, x:y=1.4:1, or x=1.4y

OpenStudy (inkyvoyd):

z:y=0.33:1 or, 0.33y=z

OpenStudy (inkyvoyd):

Now, you want to find the relation of x to z.

OpenStudy (anonymous):

...^^; Do I replace x with 0?

OpenStudy (inkyvoyd):

erm

OpenStudy (inkyvoyd):

You'll want to find y in both equations, actually

OpenStudy (anonymous):

x+z=1.73y?

OpenStudy (inkyvoyd):

erm

OpenStudy (inkyvoyd):

x=1.4y 0.33y=z x/1.4=y z/0.33=y since "x/1.4" and "z/0.33" are both equal to y, we can set them equal. z/0.33=x/1.4 Then, we simplify.

OpenStudy (anonymous):

So I mulitply .33 and 1.4 to get x?